please help me solve this word problem. It is a PUZZLE WORD PROBLEM.

Mary is three years older than her twin brothers. Next year, the sum of the ages of the three will be exactly 102. How old are they now?

Anything about this problem would be helpful. Thanks =]

Answer by fanism(39)

let Mary be M years old and her brother be B years old
Mary is three years older than her twin brothers, therefore
M = 3 + B -- eq1
Next year, the sum of the ages of the three will be exactly 102.
Since each one of them will grow 1 year old, therefore,
102 = (M+1) + (B+1) + (B+1) -- eq2
Sub eq1 into eq2,
102 = (3+B+1) + (B+1) + (B+1)
102 = 3B + 6
96 = 3B
B = 32
Put B = 32 into eq1,
M = 3 + 32
M = 35